If D=MB(Q) is P=100-5Q,S=MC(Q) is P=20+10Q.

(1)calculate Pe and Qe

(2)If Tax=20 On Seller, what is Cs Ps GS and DL.

(3)About GS,seller and Buyer which side take more Tax Burden

Please Draw graphs and answer.

 


1
Expert's answer
2021-04-15T07:15:21-0400

(1) In equilibrium D = S, so:

100 - 5Q = 20 + 10Q,

15Q = 80,

Qe = 5.33 units,

Pe = 20 + 10×5.33 = 73.33.

(2) If Tax of 20 is put on Seller, then consumer, producer and total surpluses Cs, Ps, GS will decrease, and the deadweight loss DL will occur.

(3) As the supply cirve is steeper and less elastic, then the seller will take more of the Tax Burden.